The most abundant isotope of uranium is 238U. All naturally occurring fluorine is 19F. What is the total number of neutrons contained in a molecule of UF6 formed from these isotopes

A) 146 neutrons

B) 152 neutrons

C) 200 neutrons

D) 206 neutrons

E) 352 neutrons
Answer #1

Atomic number of U = 92

mass number of U = 238

So,

number of neutron in U = mass number - atomic number

= 238 - 92

= 146

Atomic number of F = 9

mass number of F = 19

So,

number of neutron in F = mass number - atomic number

= 19 - 9

= 10

So, total number of neutrons in UF6 = number of neutron in U + 6*number of neutron in F

= 146 + 6*10

= 146 + 60

= 206

Answer: D